OpenStudy (anonymous):

How many x intercepts does the graph of the function f(x) = x 3 + 4 have?

OpenStudy (anonymous):

f(x) = 3x + 4 has one x-intercept 3x + 4 = 0 3x = -4 x = -4/3

OpenStudy (anonymous):

Is the question x^3+4?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

You'll still have 1 x intercept, but there would also be 2 imaginary roots.

OpenStudy (campbell_st):

it wll have up to 3 intercepts... so you need to solve... so if y = x^3 +4 to find the intercepts let y = 0 so x^3 + 4 = 0 solve for x that will give the intercept(s)

OpenStudy (anonymous):

There are a total of three x values, with one being a real value and two being imaginary. the real value is: x=0 and the imaginary value is x=-2i and x=2i

OpenStudy (campbell_st):

there is only 1 intercept.... \[x = \sqrt[3]{-4}\]

OpenStudy (anonymous):

yea tanx